# GenTorrent: Scaling Large Language Model Serving with An Overlay Network

### v3

@@ 4 Security Analysis and Evaluation
This section presents the analysis and evaluation on how the proposed protocols can achieve the security properties.
@@ 4.1 Anonymity
We apply a commonly used entropy-based metric [51, 34] to measure the anonymity of a system, whose definition is presented in the Appendix. It ranges from 0 and 1, and a higher value represents better anonymity.
We use simulations to compare the normalized entropy across three systems: onion routing [36, 13], garlic cast (GC) [34], and GenTorrent, in a 10,000-node network under different fractions of malicious nodes. As shown Fig. 11, GenTorrent consistently achieves the highest level of anonymity, particularly in low-to-moderate corruption scenarios. For instance, at a corruption fraction of $f=0.05$, the normalized entropy for GenTorrent reaches $0.965$, compared to $0.954$ for onion and $0.903$ for GC. While introducing path diversity enhances resilience to individual path failures, it simultaneously increases the risk of de-anonymization if adversaries are capable of cross-path collusion. However, GenTorrent uses different path IDs for different paths, hence achieving greater robustness against colluding adversaries.
@@ 4.2 Message confidentiality
The content of a prompt or response is revealed to adversaries if the colluding adversaries exist on more than $k$ paths of the user node. where $k$ is the decoding threshold for S-IDA  [37, 42, 23]. Even if adversaries control $k$ paths, exhaustive brute-force decoding over all possible combinations of forwarded messages remains computationally prohibitive.
As shown in Fig. 11, both GenTorrent and GC show degradation in confidentiality if brute-force decoding is possible – a big assumption on adversaries’ capability – with GC dropping to $0.73$ and GenTorrent to $0.88$ under high adversarial presence ($f=10\%$). If brute-force decoding is impossible, both systems maintain almost perfect confidentiality across all scenarios.
@@ 4.3 Detecting Dishonest Model Nodes
We evaluate the effectiveness and robustness of GenTorrent’s reputation mechanism using a mixture of honest and dishonest LLMs. Model nodes that do not reply will be easily detected by the verification nodes, so we focus on detecting dishonest model nodes that use low-quality models or forged responses. Note that a model node cannot differentiate the requests from normal users and those from verification nodes, because all requests are routed through the anonymous overlay.
We use Meta-Llama-3.1-8B-Instruct-Q4_0 as the ground truth (GT) model for evaluation, representing a high-quality response baseline. To simulate dishonest or degraded behavior, we introduce four alternative models with lower capabilities or altered configurations. Specifically, m1 refers to Llama-3.2-3B-Instruct-Q4_K_M, m2 refers to Llama-3.2-1B-Instruct-Q4_K_M, m3 refers to Llama-3.2-1B-Instruct-Q4_K_S, and m4 refers to Llama-3.2-3B-Instruct-Q4_K_S. These models exhibit varying levels of performance degradation due to either reduced model size or quantization settings, and serve to evaluate the sensitivity and discriminative power of the GenTorrent verification and reputation framework.
We first run 50 prompts and evaluate the responses from all five models by measuring the credit scores in the normalized perplexity. Fig. 11 shows normalized perplexity-based credibility scores. We find that statistically, the ground truth model (GT) provides higher scores than other models.
@@
Figure 14: Survival probability under churn (200 nodes/min).
Fig. 12 illustrates how the reputation scores evolve over time across 35 epochs by varying three penalty sensitivity levels: $\gamma=1$, $\frac{1}{3}$, and $\frac{1}{5}$. The results show a clear separation between GT and the weaker models, validating the effectiveness of the token-level probabilistic verification strategy. Fig. 12(a) shows the reputation trajectories under Level-1, where the penalty is relatively lenient. In this setting, the dishonest models experience slower degradation and tend to stabilize around reputation scores of 0.2 - 0.4, which may allow them to remain in the trusted pool for longer periods. With a stricter penalty in $\gamma=\frac{1}{3}$ Fig. 12(b), reputation scores of dishonest models exhibit a more significant decline, often dropping below 0.2 by period 5. For the most aggressive punishment threshold $\gamma=\frac{1}{5}$ in Fig. 12(c), results in an early and sharp decline of reputation scores, with most dishonest models falling below 0.1 within the first 5 periods.
@@ 4.4 Other Security Properties
Counterfeiting. There are three main types of counterfeiting attacks by a malicious verification leader. 1) The leader attempts to degrade some model nodes by sending them modified challenge prompts different from the pre-determined prompts agreed upon by the committee. Since model node responses always include the original prompt, other verification nodes can detect this deviation by comparing the embedded prompt in the response against the expected prompt from the pre-determined prompt list. 2) The leader alters the responses from the model nodes before broadcasting them to the committee. In GenTorrent, each response from a model node includes its digital signature, which ensures the integrity and authenticity of the responses. Any tampering by the malicious leader would be easily detected. Thus, the leader can only broadcast the original responses to the committee for scoring, but cannot alter the responses. 3) The leader can intentionally skip broadcasting responses from certain model nodes or falsely claim that a model node returned an “invalid response". To prevent this, all verification nodes reach consensus in advance on the list of model nodes to be challenged for the epoch. If verification nodes detect that some expected responses are missing or marked as invalid without justification, they will initiate their own challenge prompts to those model nodes, which are distinct from the original prompts used by the leader, to prevent auditing detection by the model nodes. If more than 2/3 of the verification nodes receive valid responses from the model nodes, the system will identify the leader as malicious. In such a case, the current epoch is aborted, and the verification process resumes in the next epoch with a newly selected leader.
@@ 5 Performance Evaluation
This section presents the performance evaluation results.
@@ 5.1 Methodology
We deploy the testbed of GenTorrent in a public cloud. The testbed includes eight model nodes with mid-tier hardware, each carrying one NVIDIA RTX A6000 GPU (48 GB) and serving a Meta-Llama-3 8B model, eight other model nodes with high-performance hardware, each carrying one NVIDIA A100 GPU (80 GB) and serving a DeepSeek-R1 14B model. We run two machines as the verification nodes. One carries an NVIDIA A100 (40 GB SXM4), 30 virtual CPUs, 200 GiB of RAM, and a 0.5 TiB SSD, and the other carries an NVIDIA GH200 (96 GB HBM), 64 virtual CPUs, 432 GiB of RAM, and 4 TiB SSD. There is no centralized coordinator in the testbed. Since we want to obtain results close to the Internet environment, we let each node delay each packet for 100ms for the Internet latency. However, we observe that the overall latency will be dominated by LLM responses, hence the Internet latency does not play a critical role. We use the most up-to-date version of vLLM [24, 45], a recently developed LLM serving library, on each single model node.
@@
1. ToolUse (ToolBench[18]): Comprising 210k queries, this workload features medium-length prompts (averaging 7,206 tokens after Zipf-1.1 sampling) with tool-specific instructions. Prefix sharing is moderate, with each query capped at a 100-token output for our tested models.
2. Coding (APPS[19]): Comprising 10,000 coding problems, this workload presents coding problems followed by detailed solution requests, averaging 1,802 tokens per prompt after Zipf-0.8 sampling. Prefix overlap is minimal, and generated outputs are longer, capped at 1,000 tokens per query for our tested models.
3. Long-Doc QA (LooGLE[27]): Comprising 776 long documents and 6.4k questions. LooGLE forms long prompts, each query includes a long document as prefix and then a question about the document, averaging 10,9852 tokens per prompts after Zipf-0.6 sampling. This scenario reflects realistic document-centric query patterns, with outputs capped at 100 tokens per query.
4. Mixed Workload: Emulating realistic heterogeneous workload conditions, this workload integrates the ToolUse, Coding, and Long-Doc QA workloads in a 3:6:1 ratio[41, 47], averaging 9,959 tokens per prompts.
@@ 5.2 Performance of Anonymous Routing
Figure 18: LLM throughput.
We evaluate the performance overhead of anonymous message (clove) processing using the Toolbench dataset [18]. Clove preparation is performed on a model node equipped with a single NVIDIA A100 40GB SXM4 GPU, 30 vCPUs, 200 GiB RAM, and 0.5 TiB SSD running Llama-3.3-70B. Decryption is executed on a user node with a single Intel i7-7700 CPU @ 3.60GHz. Figure 14 shows the cumulative distribution functions (CDFs) for both preparation and decryption latency over 10,000 trials.
@@
We conduct simulations to evaluate the communication survival rate under network churn in a 3,119-node network and a churn rate of 200 nodes/min – a very high rate. A survived communication is defined as the user successfully sending the request to a model node and decoding the response by receiving at least $k$ cloves. The results are shown in Fig. 14. The survival rates of GenTorrent and GC are both much higher than that of onion. After 15 minutes, the survival rate of GenTorrent is still above 80%, without finding new proxies. Hence, we suggest each user node discover new proxies every 15 minutes.
5.3 Performance of Overlay Forwarding
We evaluate three types of latency of the model node responding to user requests: average generation latency capped at 1000 tokens (Avg), 99th percentile (P99), and Time to First Token (TTFT). The LLMs used in the evaluations are Meta Llama 3 (8B) running on A6000 and DeepSeek-R1 (14B) running on A100. Each data point reported here is the average of three independent runs of the entire workload. Latency results are shown in Fig. 15 for GenTorrent and the design without HR-tree based KV cache reusing. GenTorrent consistently demonstrates significant latency improvements compared to the design without KV cache reusing, across all evaluated workloads, LLM, and GPU setups. Notably, under heavy workload (LooGLE and Mixed), the difference is more evident, indicating the effectiveness of GenTorrent’s overlay forwarding and cache-reuse strategy. The TTFT metric directly impacts perceived user responsiveness, a critical factor in practical deployments. GenTorrent provides significantly shorter latency in TTFT. For both LLMs, GenTorrent reduces TTFT at higher request rates by 40-50%. In addition, for higher request rates, GenTorrent maintains slower latency growth.
@@ 5.3 Performance of Overlay Forwarding
We show the KV cache hit rates of the mixed workload on 8B model nodes in Fig. 18. GenTorrent consistently achieves significantly higher cache hit rates across all evaluated workloads.Improvements in cache hit rate translate directly into throughput gains as shown in Fig. 18. We set the throughput of GenTorrent under ToolUse to 1 and show the normalized throughput. GenTorrent outperforms the design without HR-tree across all workloads.
@@ 6 Related Works
With the growing importance and popularity of LLM applications, LLM serving engines, software frameworks to run LLM inferences efficiently, have been studied and developed [24, 40, 49, 26, 41, 31, 14, 39, 46, 50, 33, 29, 28, 25]. One key challenge of an LLM serving engine is the management of KV cache. For example, vLLM [24] stores KV cache as non-continuous blocks, each with a fixed number of tokens. It enables sharing KV caches for different user requests with similar prompts.
Distributed LLM serving studies the management of a group of serving nodes in a cloud/cluster environment [17, 21, 49, 41, 29]. One key research problem is to distribute requests to different serving nodes [21, 49, 41]. SGLang [49] and Preble [41] use a radix tree to find reusable KV caches across different user requests and implement scheduling algorithms to achieve load balancing. DistServe [50] assigns prefill and decoding to different nodes to eliminate their interferences. These studies assume a centralized “scheduler” or “router” that analyzes all user requests based on various scheduling algorithms and sends the requests to serving nodes. In contrast, there is no such central scheduler in GenTorrent with knowledge of the entire system. GenTorrent is the first study of the three key problems in decentralized serving.
In the literature of Internet-based LLM serving, Petals [5] proposes to run an LLM (such as BLOOM-176B) on multiple consumer GPUs collaboratively. The Knowledge Delivery Network (KDN) [10] allows storing and exchanging KV caches among different serving engines. They focus on completely different design objectives from GenTorrent.

### v4

@@ 4.1 Anonymity
We apply a commonly used entropy-based metric [zhuang2005cashmere, garlic] to measure the anonymity of a system, whose definition is presented in the Appendix. It ranges from 0 and 1, and a higher value represents better anonymity.
We use simulations to compare the normalized entropy across three systems: onion routing [Onion1998, Tor04], garlic cast (GC) [garlic], and PlanetServe, in a 10,000-node network under different fractions of malicious nodes. As shown Fig. 10, PlanetServe consistently achieves the highest level of anonymity, particularly in low-to-moderate corruption scenarios. For instance, at a corruption fraction of $f=0.05$, the normalized entropy for PlanetServe reaches $0.965$, compared to $0.954$ for onion and $0.903$ for GC. While introducing path diversity enhances resilience to individual path failures, it simultaneously increases the risk of de-anonymization if adversaries are capable of cross-path collusion. However, PlanetServe uses different path IDs for different paths, hence achieving greater robustness against colluding adversaries.
@@ 4.2 Message confidentiality
The content of a prompt or response is revealed to adversaries if the colluding adversaries exist on more than $k$ paths of the user node. where $k$ is the decoding threshold for S-IDA  [rivest1997all, stinson2001something, SIDA93]. Even if adversaries control $k$ paths, exhaustive brute-force decoding over all possible combinations of forwarded messages remains computationally prohibitive. In Fig. 10, both PlanetServe and GC show degradation in confidentiality if brute-force decoding is possible – a big assumption on attack’s capability – with GC dropping to $0.73$ and PlanetServe to $0.88$ under high adversarial presence ($f=10\%$). If brute-force decoding is impossible, both systems maintain almost perfect confidentiality across all scenarios.
@@ 4.3 Detecting Dishonest Model Nodes
We focus on detecting dishonest model nodes that use low-quality models or forged responses. Note that a model node cannot differentiate the requests from normal users and those from verification nodes, because all requests are routed through the anonymous overlay.
We use Meta-Llama-3.1-8B-Instruct-Q4_0 as the ground truth (GT) model for evaluation, representing a reference response. To simulate dishonest or degraded behavior, we introduce four alternative models with lower capabilities or altered configurations. Specifically, m1 refers to Llama-3.2-3B-Instruct-Q4_K_M, m2 refers to Llama-3.2-1B-Instruct-Q4_K_M, m3 refers to Llama-3.2-1B-Instruct-Q4_K_S, and m4 refers to Llama-3.2-3B-Instruct-Q4_K_S. These models exhibit varying levels of performance degradation due to either reduced model size or quantization settings, and serve to evaluate the sensitivity and discriminative power of the PlanetServe verification and reputation framework. In addition, we include two further settings based on the ground-truth model but with deliberate prompt alterations. gt_cb, where prompts are rewritten into clickbait-style headlines, and gt_ic, where prompts are followed by injected long-form continuations spanning diverse genres (e.g., fictional expansions, documentary-like prose, or pop-psychology explanations).
We first run 50 prompts and evaluate the responses from all five models by measuring the credit scores in the normalized perplexity. Fig. 10 shows normalized perplexity-based credibility scores. Each point corresponds to the score of a single reply. We find that statistically, the ground truth model (GT) provides higher scores than other models. A false positive is possible, but in the long term, it will be corrected if the node follows the protocol and its credit score converges to the correct value. Fig. 11 illustrates how the reputation scores evolve over time across 35 epochs (50 prompts each) by varying three penalty sensitivity levels: $\gamma=1$, $\frac{1}{3}$, and $\frac{1}{5}$. The results show a clear separation between GT and the weaker models after the first epoch, validating the effectiveness of the token-level probabilistic verification strategy. Fig. 11(a) shows the reputation trajectories under Level-1, where the penalty is relatively lenient. In this setting, the dishonest models experience slower degradation and tend to stabilize around reputation scores of 0.2 - 0.4, which may allow them to remain in the trusted pool for longer periods. With a stricter penalty in $\gamma=\frac{1}{3}$ Fig. 11(b), reputation scores of dishonest models exhibit a more significant decline, often dropping below 0.2 by period 5. For the most aggressive punishment threshold $\gamma=\frac{1}{5}$ in Fig. 11(c), results in an early and sharp decline of reputation scores, with most dishonest models falling below 0.1 within the first 5 periods. Although the paper only shows a few set of experiments, we have run many similar experiments and PlanetServe can detect weaker models in every set of experiments.
@@
Figure 13: Survival probability and delivery rate under churn (200 nodes/min).
@@ 4.4 Other Security Properties
Counterfeiting. There are three main types of counterfeiting attacks by a malicious verification leader. 1) The leader attempts to degrade some model nodes by sending them modified challenge prompts different from the agreed prompt list. Since model node responses always include the original prompt, other verification nodes can detect this deviation. 2) The leader alters the responses from the model nodes before broadcasting them to the committee. In PlanetServe, each response from a model node includes its digital signature, which ensures the integrity and authenticity of the responses. Any tampering by the malicious leader would be easily detected. 3) The leader can intentionally skip broadcasting responses from certain model nodes or falsely claim that a model node returned an “invalid response". To prevent this, all verification nodes reach consensus in advance on the list of model nodes to be challenged for the epoch. If verification nodes detect that some expected responses are missing or marked as invalid without justification, they will initiate their own challenge prompts to those model nodes, which are distinct from the original prompts used by the leader, to prevent auditing detection by the model nodes. If more than 2/3 of the verification nodes receive valid responses from the model nodes, the system will identify the leader as malicious.
Sybil attacks. An adversary could potentially leverage Sybil attacks to control multiple verification nodes to alter verification results. Verification nodes are operated by organizations or individuals with verifiable credentials (e.g., recognized academic or research institutions). These credentials will be periodically checked and revoked if a verification node is compromised, and then the node will be expelled from the committee. Moreover, verification nodes are selected to be geographically and administratively diverse, which makes it hard for an adversary to compromise $\geq 1/3$ of the committee. To further improve the security of the verification committee, the committee members can be non-static. They are periodically rotated and re-elected. Misbehaving verification nodes will be detected and removed. This could increase the cost and difficulty for the adversary to compromise the committee.
@@ 5.1 Methodology
We deploy the testbed of PlanetServe in a public cloud. The testbed includes eight model nodes with mid-tier hardware, each carrying one NVIDIA RTX A6000 GPU (48 GB) and serving a Meta-Llama-3 8B model, eight other model nodes with high-performance hardware, each carrying one NVIDIA A100 GPU (80 GB) and serving a DeepSeek-R1-Qwen-14B model. We run two machines as the verification nodes. One carries an NVIDIA A100 (40 GB SXM4), 30 virtual CPUs, 200 GiB of RAM, and a 0.5 TiB SSD, and the other carries an NVIDIA GH200 (96 GB HBM), 64 virtual CPUs, 432 GiB of RAM, and 4 TiB SSD. There is no centralized coordinator in the testbed. Internet latency in simulations is from samples collected by our measurement. We use the most up-to-date version of vLLM [vllm, vllmgithub], a recently developed LLM serving library, on each single model node.
@@
1. ToolUse (ToolBench[guo2024stabletoolbench]): .210k queries with tool-specific instructions; prompts average 7,206 tokens after Zipf-1.1 sampling. Prefix sharing is moderate, and outputs are capped at 100 tokens per query for the models we evaluate.
2. Coding (APPS[hendrycksapps2021]): .10,000 coding problems with detailed solution requests; prompts average 1,802 tokens after Zipf-0.8 sampling. Prefix overlap is minimal, and outputs are capped at 1,000 tokens per query for the models we evaluate.
3. Long-Doc QA (LooGLE[li2023loogle]): 776 long documents paired with 6.4k questions. Each prompt includes a long document as prefix followed by a question, averaging 10,985 tokens per prompt after Zipf-0.6 sampling. Outputs are capped at 100 tokens per query for the models we evaluate.
4. Mixed Workload: It integrates the ToolUse, Coding, and Long-Doc QA workloads in a 3:6:1 ratio per real-world traces [Preble25, wu2023fast], averaging 9,959 tokens per prompts (All token counts above are computed using the Llama 3 tokenizer).
@@ 5.2 Anonymous Routing
Figure 17: Norm. LLM throughput
We evaluate the performance overhead of anonymous message (clove) processing using the Toolbench dataset [guo2024stabletoolbench]. Clove preparation is performed on a model node equipped with a single NVIDIA A100 40GB SXM4 GPU, 30 vCPUs, 200 GiB RAM, and 0.5 TiB SSD running Llama-3.3-70B. Decryption is executed on a user node with a single Intel i7-7700 CPU @ 3.60GHz. Figure 13 shows the cumulative distribution functions (CDFs) for both preparation and decryption latency over 10,000 trials.
@@
We conduct simulations to evaluate the communication survival rate under network churn in a 3,119-node network and a churn rate of 200 nodes/min – a very high rate. We deployed 14 AWS t3.micro instances across multiple regions and measured end-to-end routing latency with realistic prompt payloads over 4,000 runs (3 nodes randomly fail in the USA; 1 randomly fails worldwide). We incorporate latency, link failures, packet loss, and congestion to the simulation and show the results in Fig. 13. The path survival rate of PlanetServe is the highest among these methods. When there is no failure, all methods have high delivery rates. However, when failure occurs, only PlanetServe can maintain a similar high delivery rate. The delivery rate of Onion significantly decreases.
5.3 Latency of Confidential Computing (CC)
model mean P99
$C$-on $C$-off $C$-on $C$-off
Llama-3.1 8B 132.19 130.95 123.94 122.70
DS-R1-Q 14B 211.58 210.96 197.9527 195.77
Table 1: Latency comparison under CC mode.
In Table 1, we show the results of the experiments on an Azure Standard NCC40ads H100 v5 virtual machine (Confidential VM), equipped with a single NVIDIA H100 GPU, 40 vCPUs, and 320 GB RAM, for the CC-on setting. For the CC-off setting, we used the corresponding Standard NC40ads H100 v5 VM, which provides the same hardware resources but does not enable confidential computing features. In both settings, the request rate was fixed at 20 requests per second. The results show that CC adds very small latency.
5.4 Overlay Forwarding
We evaluate three types of latency of the model node responding to user requests: average generation latency capped at 1000 tokens (Avg), 99th percentile (P99), and Time to First Token (TTFT). The LLMs used in the evaluations are Meta Llama 3 (8B) running on A6000 and DeepSeek-R1 (14B) running on A100. We compare PlanetServe with a centralized baseline of 8 A100 GPUs without KV-cache sharing. Each data point reported here is the average of three independent runs of the entire workload. Latency results are shown in Fig. 14 for PlanetServe and the centralized baseline. PlanetServe consistently demonstrates significant latency improvements across all evaluated workloads, LLM, and GPU setups. Notably, under heavy workload (LooGLE and Mixed), the difference is more evident, indicating the effectiveness of PlanetServe’s overlay forwarding and cache-reuse strategy. The TTFT metric directly impacts perceived user responsiveness, a critical factor in practical deployments. PlanetServe provides significantly shorter latency in TTFT. For both LLMs, PlanetServe reduces TTFT at higher request rates by 40-50%. In addition, for higher request rates, PlanetServe maintains slower latency growth.
@@ 5.4 Overlay Forwarding
We show the KV cache hit rates of the mixed workload on DeepSeek-R1-Qwen-14B model nodes in Fig. 17. In addition to the centralized baseline, we further compare PlanetServe to centralized serving with 8× A100 GPUs, running vLLM with the default continuous-batching scheduler and tensor parallelism. PlanetServe consistently achieves significantly higher cache hit rates across all evaluated workloads. The throughput comparison is shown in Fig. 17. PlanetServe outperforms the baseline without sharing, and tensor parallelism provides the highest throughput due to the scheduler and parallelism.
@@ 5.5 Verification throughput
6 Discussion
Lesson learned from past projects. We investigate the challenges of the past decentralized overlay networks, such as PlanetLab and P2P file sharing, and discuss how the design of PlanetServe can learn from these projects. 1) Incentive design is important. Performance efficiency alone does not necessarily lead to adoption. Hence, we design a comprehensive incentive infrastructure with verification services. We plan to keep improving this infrastructure. 2) Churn-resilience and security should be considered at the beginning of deployment, which are the main research objectives of this work. 3) Integration with modern computing architecture. Some past P2P networks cannot be integrated with the cloud architecture, which significantly limits their applications. The design of PlanetServe aims at a flexible and scalable architecture that is compatible with cloud and can evolve with new technologies.
More complex models. This paper focuses on text-only tasks. Future research will extend the architecture to support more complex models. A key direction is serving multimodal models including requests with non-textual data. In fact, HR-Tree provides advantages in supporting non-textual data compared to existing KV cache sharing methods, because it records the hash values rather than the specific texts. Another future research direction is to support advanced reasoning and tool-using models. The workloads of these models could be different from the current ones, requiring adaptation of PlanetServe’s current KV cache sharing and load-balancing mechanisms. However, the main infrastructure of PlanetServe can remain the same. New optimization methods can be deployed in it incrementally.
Model partitioning. Due to the limited size of a research paper, we do not consider model partitioning [borzunov2023petal], which can be complementary to PlanetServe. The new combined research objective, i.e., how to achieve decentralized overlay LLM serving with model partitioning, can be future research.
Blockchain integration. Our committee-based BFT mechanism provides a lightweight and efficient solution for serving and verification without incurring the additional overhead of deploying blockchain. However, blockchain offers desirable properties such as transparency, auditability, and simplified global coordination. In particular, smart contracts could automate reputation management in a transparent and tamper-resistant manner. Thus, We will integrate blockchain in the future to enhance transparency and auditability as the system scales to support a larger number of users and model nodes.
@@ 7 Related Works
With the growing importance and popularity of LLM applications, LLM serving engines, software frameworks to run LLM inferences efficiently, have been studied and developed [vllm, PowerInfer, sglang24, Speculative, Preble25, TensorRT, TGI, Fairness2024Sheng, LoongServe2024, DistServe2024, SplitWise2024, CacheGen2024, liu2024droidspeak, InfiniGen]. One key challenge of an LLM serving engine is the management of KV cache. For example, vLLM [vllm] stores KV cache as non-continuous blocks, each with a fixed number of tokens. It enables sharing KV caches for different user requests with similar prompts.
Distributed LLM serving studies the management of a group of serving nodes in a cloud/cluster environment [ServerlessLLM, IntelligentRouter, sglang24, Preble25, CacheGen2024]. One key research problem is to distribute requests to different serving nodes [IntelligentRouter, sglang24, Preble25]. SGLang [sglang24] and Preble [Preble25] use a radix tree to find reusable KV caches across different user requests and implement scheduling algorithms to achieve load balancing. DistServe [DistServe2024] assigns prefill and decoding to different nodes to eliminate their interferences. These studies assume a centralized “scheduler” or “router” that analyzes all user requests based on various scheduling algorithms and sends the requests to serving nodes. In contrast, there is no such central scheduler in PlanetServe with knowledge of the entire system. PlanetServe is the first study of the three key problems in decentralized serving.
Petals [borzunov2023petal] proposes to run an LLM (such as BLOOM-176B) on multiple consumer GPUs collaboratively. The Knowledge Delivery Network (KDN) [kdn] allows storing and exchanging KV caches among different serving engines. DiLoCo [diloco2024] uses distributed clusters for training. Orca [Orca2022] uses a centralized scheduling algorithm to improve serving latency and throughput. None of them address the challenges discused in PlanetServe, including decentralized serving and verification, overlay forwarding, and anonymous communication.